# Leveraging Predictive Equivalence in Decision Trees

### v1

@@ 1 Introduction
(1) Decision trees imply an evaluation procedure that can get stuck on irrelevant missing information. If $x_{1}$ is missing and $x_{2}=0$, the first tree in Figure 1 cannot be traversed, but the second tree clearly predicts 00.
@@ 2.1 Decision Trees and Simplicity
Our work is particularly related to the problem of explanation redundancy in decision trees. This concept is explored by Izza et al. (2022), who demonstrate that the paths taken through the tree to reach predictions (“path explanations”) often have redundant variables in them, which are not necessary to make the prediction. The authors present polynomial-time algorithms to compute succinct path explanations. In contrast, we present a method to compute a minimal boolean logical representation of the entire decision tree, using the Quine-McCluskey algorithm (Quine, 1952; McCluskey, 1956) as a subroutine. This representation enables succinct path explanations of predictions for free, once the global representation is computed for some up-front cost. Our representation also enables several downstream applications beyond prediction explanations.
A line of work on the simplicity of machine learning models shows that when data has noise in the outcomes (common on many tasks we consider in our experiments), simpler decision trees will be competitive in performance with more complicated ones (Semenova et al., 2022, 2023; Boner et al., 2024). If our decision trees have a small number of leaves, the number of variables in the Quine-McCluskey subroutine will be small, and our algorithm for simplification will be efficient despite the NP-completeness of the problem.
@@ 2.2 Applications
We optimize the cost of applying a decision tree by applying Q-learning (Watkins, 1989). Q-learning is a model-free approach for policy learning that estimates the value of each action in each state by allowing an agent to explore the state space. During exploration, the reward of the $j$-th visited state is gradually propagated back to the $(j-1)$-th state. Given sufficient episodes – iterations of this exploration – it has been shown that Q-learning will produce the optimal policy for a given problem (Watkins, 1989; Watkins & Dayan, 1992). Since the introduction of Q-learning, the field of reinforcement learning has dramatically expanded. We refer readers to recent survey papers on the field for a more complete literature review (Shakya et al., 2023; Wang et al., 2022). However, we found that the update rule and regime proposed by (Watkins, 1989) were sufficient.
@@ 3 Methodology
Consider a dataset $D=\{(x_{i},y_{i})\}_{i=1}^{n}$, where each $x_{i}\in\mathbb{R}^{d}$ and $y_{i}\in\{0,1\}$ pair is sampled i.i.d. from some unknown distribution $\mathcal{D}$. For the purposes of our work, these $x_{i}$’s may be continuous, ordinal, categorical, or binary. We refer to the $j^{th}$ feature of the $i^{th}$ sample as $x_{i,j}$. We use the notation $x_{\cdot,j}$ to refer to the $j^{th}$ feature. We consider binary classification problems in this work, though our results can be extended to multiclass classification with minor adjustments to the algorithms and theorems. We also work with binarized datasets, in which feature $j$ of $D$ is binarized into $B_{j}$ different binary features. For example, the feature $age$ may be binarized into binary features $age\leq 5$, $age\leq 10$, etc. We denote the $k^{th}$ binary feature corresponding to feature $j$ of the $i^{th}$ sample as $b_{i,j}^{(k)}$. We reserve capital letters $X$ and $Y$ for random variables, and we index random variables via subscripts, i.e., $X_{i}$.
@@ 3.1 Representing Trees to Resolve Predictive Equivalence
{forest}
[$X_{1}$, [$X_{3}$ [0] [1] ] [$X_{2}$ [0] [1] ] ]
DNF $(X_{1}\wedge X_{2})\vee(\neg X_{1}\wedge X_{3})$ BCF $(X_{1}\wedge X_{2})\vee(\neg X_{1}\wedge X_{3})\vee(X_{2}\wedge X_{3})$
@@
Algorithm 1 describes how we transform trees into minimal DNF representation. This algorithm combines the positive-predicting leaves of the decision tree into an expression in disjunctive normal form. This expression is then simplified with a slightly modified version of the Quine-McCluskey algorithm (Quine, 1952) (see Algorithm 5) to find the minimal form of the boolean expression encoding positive predictions by the tree. We perform the same procedure on the negative-predicting leaves to obtain a minimal boolean expression for evaluating whether the tree predicts negative. Algorithm 2 explains how this method provides predictions, with ‘substitute’ meaning each variable with a known value is replaced by a constant (e.g., if $x_{i,1}=1$, $(x_{\cdot,1}\land x_{\cdot,2})\lor(\lnot x_{\cdot,2})$ becomes $(1\land x_{\cdot,2})\lor(\lnot x_{\cdot,2})=True$). Equivalence is defined in Algorithm 3 in Appendix B.
@@
  $SimplePosExpr\leftarrow QuineMcCluskey(PosExpr)$
  $SimpleNegExpr\leftarrow QuineMcCluskey(NegExpr)$
@@
[…]
  $\textrm{expr}\leftarrow QuineMcCluskey(expr)$
@@ 3.2 Datasets
We consider four datasets throughout this work and eight additional datasets in Appendix C. We refer to the primary four as COMPAS (Larson et al., 2016), Wine Quality (Cortez et al., 2009), Wisconsin (Street et al., 1993), and Coupon (Wang et al., 2017). COMPAS measures 7 features for 6,907 individuals, where labels are whether the individuals were arrested within 2 years of being released from prison. Wine Quality reports 11 features over 6,497 wines along with a numerical quality rating between 1 and 10. We binarize these ratings into high ($>5$) and low ($\leq 5$) quality classes and predict this binary rating. Wisconsin is a breast cancer dataset and contains 30 features over 569 masses, where labels designate whether the tumor was malignant or benign. Coupon measures 25 features for 12,684 individuals, and labels denote whether or not the individual would accept a coupon. See Appendix D.1 for complete details on the preprocessing applied to each dataset.
@@ 5.1 Gini Importance
[…]
@@ 6 Case Study 2: Missing Data
path 2.4×2.4$\times$2.4 × 3.9×3.9$\times$3.9 × 2.6×2.6$\times$2.6 × 3.4×3.4$\times$3.4 ×
features 32.8×32.8$\times$32.8 × 64.6×64.6$\times$64.6 × 28.6×28.6$\times$28.6 × 16.1×16.1$\times$16.1 ×
@@ 7 Case Study 3: Improving Cost Efficiency
Reward Function. When a feature $b_{i,j}^{(k)}$ (the $k^{th}$ bin on the $j^{th}$ feature of the $i^{th}$ example) is measured, a cost of $c_{j}$ (i.e., a reward of $-c_{j}$) is incurred if there is no $k^{\prime}$ such that $b_{i,j}^{(k^{\prime})}$ has been purchased; otherwise, no cost is incurred to reflect the fact that a practitioner would obtain the value of the input feature, not an individual bin. If enough features are known to satisfy any clause of the Blake canonical form of a tree, a reward of $\sum_{j=1}^{d}c_{j}$ is given, and the current episode of Q-learning is terminated.
@@
We evaluate the cost savings of this approach using the COMPAS, Wine Quality, Wisconsin, and Coupon datasets. For each dataset considered, we randomly generate an integer cost between $1$ and $10$ for each feature. We consider three purchasing policies: 1) following the BCF/Q-learning policy as outlined above, 2) purchasing features in the order suggested by traversing the tree, and 3) directly purchasing every feature in the tree.We then evaluate the average cost incurred by each policy across samples from the test dataset. We fit $50$ decision trees on distinct bootstrap samples of each dataset, and perform this evaluation for each tree produced.
@@ Appendix A Proofs
Case 2: $\mathcal{T}_{\textrm{DNF}}(m(x_{i}))\neq NA$. Suppose WLOG $\mathcal{T}_{\textrm{DNF}}(m(x))=1$. Then we know that the known variable assignments in $m(x)$ satisfy $SimplePoxExpr$. Therefore, for any completion $z$ of $m(x)$, $SimplePosExpr$ is satisfied by $z$. Thus, by Lemma A.1, we have $\mathcal{T}(x)=1$ for all completions of $m(x)$. ∎
@@
Let $x\in\mathbb{R}^{d}$. Assume $\mathcal{T}(x)=\hat{y}\in\{0,1\}$, and let the explanation for $\mathcal{T}_{\textrm{DNF}}(x)$ be any term in $SimplePosExpr$ satisfied by $x$ if $\hat{y}=1$, and any term in $SimpleNegExpr$ satisfied by $x$ if $\hat{y}=0$. If any term in either expression is satisfied, no variable in the explanation is redundant.
@@
By Lemma A.1, we know that $\mathcal{T}_{\textrm{DNF}}(m(x))=\hat{y}\in\{0,1\}$ if and only if $T(z)=\hat{y}$, for all completions $z$ of $m(x)$. The set of all completions of $m(x)$ includes the original values of $x$, in addition to any possible imputation of $m(x)$. Therefore, for all imputations $z$ of $m(x)$, $\mathcal{T}(z)=\mathcal{T}(x)$. ∎
@@ Appendix B Algorithm Details
Algorithm 5 Adjusted Quine-McCluskey (with particular processing of the data structures involved to allow for proof of Theorem 3.4)
@@ C.3 Additional Results for Case Study 1: Variable Importance
Netherlands $>$20 previous case 0.220336 0.000000
@@ D.2 Hyperparameter Settings
In all of our experiments with RID, we computed RID over 100 bootstrap iterations with an additive Rashomon bound of 0.02, a sparsity penalty of 0.02, and a maximum tree depth of 3 (i.e., the $depth\_bound$ parameter was set to 4). Note that, when using RID, we applied a larger sparsity penalty than we did in our other experiments (0.02 instead of 0.01) because we ran into computational constraints when running RID over very large Rashomon sets. We did not include “trivial extension” in these Rashomon sets because that is the default behavior of RID.
In our missing data experiments, we fit decision trees using SKLearn’s Decision Tree implementation (Pedregosa et al., 2011), with a maximum depth of 3 and all other parameters left as their default values. Our Rashomon sets used depth 3 (i.e., the $depth\_bound$ parameter was set to 4), additive Rashomon bound 0.02, and sparsity penalty 0.01. (we did not include trivial extensions because those are irrelevant to what we were investigating; a trivial extension never allows handling of additional missing data). Missingness is injected into complete datasets synthetically. For each sample, we introduced missingness independently into each binarized feature with probability $p\in[0.1,0.2,\ldots,0.9]$. We investigate alternative depths of sklearn and pre-binarization missingness in Appendix Section E.
@@ Appendix G Detailed Description of Q-learner Initialization
  function initialize($\mathcal{T}_{cur}$, $D_{cur}$, $state$
@@
        $state_{\ell}\leftarrow$ copy of $state$ with feature $j$ set to 0
        $r_{\ell}\leftarrow$ initialize$(\mathcal{T}_{cur}.left\_subtree,D_{\ell},state_{\ell})$
@@
        $state_{r}\leftarrow$ copy of $state$ with feature $j$ set to 1
        $r_{r}\leftarrow$ initialize$(\mathcal{T}_{cur}.right\_subtree,D_{r},state_{r})$
@@
  initialize($\mathcal{T}$, $D$, $state$)

### v3

@@ 1 Introduction
(1) Decision trees imply an evaluation procedure that can get stuck on irrelevant missing information. If $x_{1}$ is missing and $x_{2}=0$, the first tree in Figure 1 cannot be traversed, but the second tree clearly predicts 0.
@@ 2.1 Decision Trees and Simplicity
Our work is particularly related to the problem of explanation redundancy in decision trees. This concept is explored by Izza et al. (2022), who demonstrate that the paths taken through the tree to reach predictions (“path explanations”) often have redundant variables in them, which are not necessary to make the prediction. The authors present polynomial-time algorithms to compute succinct path explanations. In contrast, we present a method to compute a minimal boolean logical representation of the entire decision tree, using a modified Quine-McCluskey algorithm (Quine, 1952; McCluskey, 1956) as a subroutine. This representation enables succinct path explanations of predictions for free, once the global representation is computed for some up-front cost. Our representation also enables several downstream applications beyond prediction explanations.
A line of work on the simplicity of machine learning models shows that when data has noise in the outcomes (common on many tasks we consider in our experiments), simpler decision trees will be competitive in performance with more complicated ones (Semenova et al., 2022, 2023; Boner et al., 2024). If our decision trees have a small number of leaves, the number of variables in the Quine-McCluskey subroutine will be small, and our algorithm for simplification will be efficient despite the NP-hardness of finding a minimal DNF expression.
@@ 2.2 Applications
We optimize the cost of applying a decision tree by applying Q-learning (Watkins, 1989). Q-learning is a model-free approach for policy learning that estimates the value of each action in each state by allowing an agent to explore the state space. During exploration, the reward of the $j$-th visited state is gradually propagated back to the (j1)(j-1)-th state. Given sufficient episodes – iterations of this exploration – it has been shown that Q-learning will produce the optimal policy for a given problem (Watkins, 1989; Watkins & Dayan, 1992). Since the introduction of Q-learning, the field of reinforcement learning has dramatically expanded. We refer readers to recent survey papers on the field for a more complete literature review (Shakya et al., 2023; Wang et al., 2022). However, we found that the update rule and regime proposed by (Watkins, 1989) were sufficient.
@@ 3 Methodology
Consider a dataset $D=\{(x_{i},y_{i})\}_{i=1}^{n}$, where each $x_{i}\in\mathbb{R}^{d}$ and $y_{i}\in\{0,1\}$ pair is sampled i.i.d. from some unknown distribution $\mathcal{D}$. For the purposes of our work, these $x_{i}$’s may be continuous, ordinal, categorical, or binary. We refer to the $j^{th}$ feature of the $i^{th}$ sample as $x_{i,j}$. We use the notation $x_{\cdot,j}$ to refer to the $j^{th}$ feature. We consider binary classification problems in this work, though our results can be extended to multiclass classification with minor adjustments to the algorithms and theorems. We also work with binarized datasets, in which feature $j$ of $D$ is binarized into $B_{j}$ different binary features. For example, the feature ageage may be binarized into binary features $age\leq 5$, $age\leq 10$, etc. We denote the $k^{th}$ binary feature corresponding to feature $j$ of the $i^{th}$ sample as $b_{i,j}^{(k)}$. We reserve capital letters $X$ and $Y$ for random variables, and we index random variables via subscripts, i.e., $X_{i}$.
@@ 3.1 Representing Trees to Resolve Predictive Equivalence
[…]
DNF $(X_{1}\wedge X_{2})\vee(\neg X_{1}\wedge X_{3})$
BCF $(X_{1}\wedge X_{2})\vee(\neg X_{1}\wedge X_{3})\vee(X_{2}\wedge X_{3})$
@@
Algorithm 1 describes how we transform trees into minimal DNF representation. This algorithm combines the positive-predicting leaves of the decision tree into an expression in disjunctive normal form. This expression is then simplified with a slightly modified version of the Quine-McCluskey algorithm (Quine, 1952) (see Algorithm 5) to find a minimal form of the boolean expression encoding positive predictions by the tree. It is important to note that there is no canonical minimal representation of arbitrary boolean expressions; the modifications in Algorithm 5 ensure that the procedure outputs the same minimal representation for equivalent expressions. We perform the same procedure on the negative-predicting leaves to obtain a minimal boolean expression for evaluating whether the tree predicts negative. Algorithm 2 explains how this method provides predictions, with ‘substitute’ meaning each variable with a known value is replaced by a constant (e.g., if $x_{i,1}=1$, $(x_{\cdot,1}\land x_{\cdot,2})\lor(\lnot x_{\cdot,2})$ becomes $(1\land x_{\cdot,2})\lor(\lnot x_{\cdot,2})=True$). Equivalence is defined in Algorithm 3 in Appendix B.
@@
$SimplePosExpr\leftarrow\textrm{QuineMcCluskey${}^{*}$}(PosExpr)$
$SimpleNegExpr\leftarrow\textrm{QuineMcCluskey${}^{*}$}(NegExpr)$
@@
$\textrm{expr}\leftarrow$ Substitute known feature values of $m(x)$ into $\mathcal{T}_{\textrm{DNF}}.SimplePosExpr$
$\textrm{expr}\leftarrow\textrm{QuineMcCluskey}(expr)$
@@ 3.2 Datasets
We consider four datasets throughout this work and eight additional datasets in Appendix C. We refer to the primary four as COMPAS (Larson et al., 2016), Wine Quality (Cortez et al., 2009), Wisconsin (Street et al., 1993), and Coupon (Wang et al., 2017). COMPAS measures 7 features for 6,907 individuals, where labels are whether the individuals were arrested within 2 years of being released from prison. Wine Quality reports 11 features over 6,497 wines along with a numerical quality rating between 1 and 10. We binarize these ratings into high (>5>5) and low ($\leq 5$) quality classes and predict this binary rating. Wisconsin is a breast cancer dataset and contains 30 features over 569 masses, where labels designate whether the tumor was malignant or benign. Coupon measures 25 features for 12,684 individuals, and labels denote whether or not the individual would accept a coupon. See Appendix D.1 for complete details on the preprocessing applied to each dataset.
@@ 5.1 Gini Importance
Predictive equivalence poses an immediate challenge for variable importance methods. To demonstrate this, we consider the toy setting where $Y=X_{1}X_{2}$ and $X_{1},X_{2}\overset{\text{i.i.d.}}{\sim}Bernoulli(0.5).$ Figure 3 presents two distinct decision trees that perfectly match this data generating process. Even in this simple case, we observe that equivalent trees can produce dramatically different variable importances when computing an impurity-based variable importance such as Gini importance. The first tree claims $X_{0}$ is twice as important as $X_{1}$ and the second tree claims the opposite.
Figure 3: Two equivalent decision trees for the setting where $Y=X_{1}X_{2}$ and $X_{1},X_{2}\overset{\text{i.i.d.}}{\sim}Bernoulli(0.5)$. Although they always produce identical predictions, achieve the same objective value, and are produced by the same algorithm, these two trees produce dramatically different variable importance values. gini refers to the Gini coefficient of each leaf, samples to the number of points falling into each leaf, and value denotes the number of negative (left) and positive (right) samples in the leaf.
This effect becomes more pronounced with more variables. We next consider a similar data generating process with $Y=\prod_{i=1}^{10}X_{i}$ and $X_{1},\ldots,X_{12}\overset{\text{i.i.d.}}{\sim}{Bernoulli(0.5)}.$ There are 12 input variables but only variables 1 though 10 are used in the data generation process, meaning there are 2 unimportant variables. We greedily fit 3 predictively equivalent decision trees over the same data using different random seeds and measured the Gini importance of each variable to each tree. Figure 4 shows the distribution of importance for each variable over these trees. We observe that the importance of each variable varies widely over predictively equivalent trees. Moreover, the importance of some useful variables is nearly indistinguishable from the importance of the extraneous variables – e.g., $X_{10}$ has importance close to 0.
@@ 6 Case Study 2: Missing Data
path $2.4\times$ $3.9\times$ $2.6\times$ $3.4\times$
features $32.8\times$ $64.6\times$ $28.6\times$ $16.1\times$
@@ 7 Case Study 3: Improving Cost Efficiency
Reward Function. When a feature $b_{i,j}^{(k)}$ (the $k^{th}$ bin on the $j^{th}$ feature of the $i^{th}$ example) is measured, a cost of $c_{j}$ (i.e., a reward of $c_{j}$) is incurred if there is no $k^{\prime}$ such that $b_{i,j}^{(k^{\prime})}$ has been purchased; otherwise, no cost is incurred to reflect the fact that a practitioner would obtain the value of the input feature, not an individual bin. If enough features are known to satisfy any clause of the Blake canonical form of a tree, a reward of $\sum_{j=1}^{d}c_{j}$ is given, and the current episode of Q-learning is terminated.
@@
We evaluate the cost savings of this approach using the COMPAS, Wine Quality, Wisconsin, and Coupon datasets. For each dataset considered, we randomly generate an integer cost between 11 and $10$ for each feature. We consider three purchasing policies: 1) following the BCF/Q-learning policy as outlined above, 2) purchasing features in the order suggested by traversing the tree, and 3) directly purchasing every feature in the tree.We then evaluate the average cost incurred by each policy across samples from the test dataset. We fit $50$ decision trees on distinct bootstrap samples of each dataset, and perform this evaluation for each tree produced.
@@ Appendix A Proofs
Case 2: $\mathcal{T}_{\textrm{DNF}}(m(x_{i}))\neq NA$. Suppose WLOG $\mathcal{T}_{\textrm{DNF}}(m(x))=1$. Then we know that the known variable assignments in $m(x)$ satisfy $SimplePoxExpr$. Therefore, for any completion $z$ of $m(x)$, SimplePosExprSimplePosExpr is satisfied by $z$. Thus, by Lemma A.1, we have $\mathcal{T}(x)=1$ for all completions of $m(x)$. ∎
@@
[…]
@@
[…]
@@ Appendix B Algorithm Details
Algorithm 5 Quine-McCluskey (Adjusted Quine-McCluskey, with particular processing of the data structures involved to allow for proof of Theorem 3.4)
@@ C.3 Additional Results for Case Study 1: Variable Importance
Netherlands >>20 previous case 0.220336 0.000000
@@ D.2 Hyperparameter Settings
In all of our experiments with RID, we computed RID over 100 bootstrap iterations with an additive Rashomon bound of 0.02, a sparsity penalty of 0.02, and a maximum tree depth of 3 (i.e., the depth_bounddepth\_bound parameter was set to 4). Note that, when using RID, we applied a larger sparsity penalty than we did in our other experiments (0.02 instead of 0.01) because we ran into computational constraints when running RID over very large Rashomon sets. We did not include “trivial extension” in these Rashomon sets because that is the default behavior of RID.
In our missing data experiments, we fit decision trees using SKLearn’s Decision Tree implementation (Pedregosa et al., 2011), with a maximum depth of 3 and all other parameters left as their default values. Our Rashomon sets used depth 3 (i.e., the depth_bounddepth\_bound parameter was set to 4), additive Rashomon bound 0.02, and sparsity penalty 0.01. (we did not include trivial extensions because those are irrelevant to what we were investigating; a trivial extension never allows handling of additional missing data). Missingness is injected into complete datasets synthetically. For each sample, we introduced missingness independently into each binarized feature with probability $p\in[0.1,0.2,\ldots,0.9]$. We investigate alternative depths of sklearn and pre-binarization missingness in Appendix Section E.
@@
D.3 Implementation Details
Our publicly available code, at https://github.com/HaydenMcT/predictive-equivalence, provides an implementation of our approach as well as further details on the experiments. In our implementation, we used the sympy library’s implementation of Quine-McCluskey simplification. We checked its suitability in place of Algorithm 5, and its satisfaction of Theorem 3.4 for ensuring correct predictive equivalence, by inspecting the source code and testing it against an exhaustive truth table checker.
@@ E.3 Results at different depths
E.4 Accelerated Missing Data Solver
Our objective in this paper was primarily to study predictive equivalence, and how DNF representations may be used to account for and leverage predictive equivalence. In doing so, we showed that the DNF representation effectively handles missing data for sparse decision trees. However, that approach requires solving a problem that is NP-hard in the number of leaves in the tree, which may not be tractable for large trees. Here, we introduce an algorithm that handles missing data scalably for larger trees. Algorithm 6 determines whether any predictively equivalent form of a given tree can be evaluated for a given sample, and if so what that tree’s prediction will be. It runs in time linear in the number of leaves of the given tree.
Algorithm 6 Missingness Evaluation
Input: $\mathcal{T}^{(1)}$, a decision tree, and $m(x)$, a sample that may have missing values
Output: $\mathcal{T}_{\textrm{DNF}}(m(x))$. (That is, NA if and only if no predictively equivalent form of $\mathcal{T}^{1}$ can be evaluated on $m(x)$. Else, returns a single prediction corresponding to any possible completion of $m(x)$)
$\textrm{nodes}\leftarrow[\mathcal{T}^{(1)}.\textrm{root}]$
 prediction $\leftarrow$ None
while $len(\textrm{nodes})>0$ do
  cur_node $\leftarrow$ nodes.pop()
  if cur_node is a leaf then
   if prediction $\neq$ cur_node.prediction and prediction $\neq$ None then
    Return NA {since two reachable leaves’ predictions do not match}
   else
    prediction $\leftarrow$ cur_node.prediction
   end if
  end if
  $f_{\textrm{cur\_node}}(m(x))$ $\leftarrow$ value of $m(x)$ at feature queried by cur_node
  if $f_{\textrm{cur\_node}}(m(x))=$ NA then
   nodes.push(cur_node.true, cur_node.false) {both children are reachable with different completions of m(x)m(x)}
  else
   if $f_{\textrm{cur\_node}}(m(x))=$ true then
    nodes.push(cur_node.true)
   else
    nodes.push(cur_node.false)
   end if
  end if
end while
 Return prediction
There exists a previous polynomial time approach for deciding weak abductive explanations in decision trees with complete assignments, but not ones with missing data (Algorithm 2 in (Izza et al., 2022)). Since the problem of evaluating a decision tree with missing data is similar to deciding a weak abductive explanation, such an approach could also be applied here if specialized to account for missing values.
@@ Appendix G Detailed Description of Q-learner Initialization
function initialize($\mathcal{T}_{cur}$, $D_{cur}$, statestate)
@@
   $state_{\ell}\leftarrow$ copy of statestate with feature $j$ set to 0
[…]
@@
   $state_{r}\leftarrow$ copy of statestate with feature $j$ set to 1
[…]
@@
 initialize($\mathcal{T}$, $D$, statestate)